# Exploring hidden priors when interpreting gravitational wave and electromagnetic probes of the nuclear equation of state

### v2

@@ II.4 Gravitational wave observations
where $x=(m_{1},m_{2},\mathbf{\chi}_{1},\mathbf{\chi}_{2})$ are binary parameters omitting tides and $X$ supplements these parameters with the EoS-derived tidal parameters (i.e., $X=({\cal M}_{c},\eta,\mathbf{\chi}_{1},\mathbf{\chi}_{2},\Lambda(m_{1}),\Lambda(m_{2}))$). For simplicity, we also use RIFT to perform this integral [122, 123], using the RIFT marginal likelihood estimator $\hat{L}(X)$ for $L(X)$. These EoS evidence infrastructure has been validated in previous work [123]. We do not employ any customary approximations when evaluating this integral numerically (e.g., that the chirp mass is well known).
While other objects putatively containing matter have been identified via their GW signature (i.e., GW190425 [134], GW200105, and GW200115 [135]), their GW signature also does not meaningfully contain the nature of nuclear matter. For example, the two objects in NSBH binaries have putative matter effects that minimally impact $\tilde{\Lambda}$. Conversely, in GW190425, the large NS masses imply very small tidal deformabilities, but that low-amplitude GW measurement only weakly constraints $\tilde{\Lambda}$. For this reason, we omit further discussion of these three objects.
Phenomenology of $\Lambda,~{}R$
To provide a preliminary estimate of the impact of the hidden NS population priors on EoS inference derived from GW170817, we follow Zhao and Lattimer 2018 [136], and adopt simple approximate relationships between tidal deformability and fiducial NS radius $R_{1.4}$. In brief, they argue that to a good approximation $\Lambda\propto(R/M)^{6}$ (Eq. (13) in Ref. [136]); that therefore $\tilde{\Lambda}$ is to a good approximation proportional to $(R_{1.4}/{\cal M}_{c})^{6}$, with a nearly mass-ratio-independent constant (Eq. (15) in Ref. [136]); and therefore for a binary with well-measured chirp mass, the radius of a fiducial NS can be estimated as (evaluated for GW170817)
$R_{1.4}\simeq 13.4~{}{\rm km}(\tilde{\Lambda}/800)^{1/6}\;.$ (10)
II.5 Additional sources of external information
The nuclear EoS must be consistent with properties of nuclei [98, 73]. Following [45], one strategy for incorporating information about nuclear-density matter is to require the baryon density $n$ and energy density $\rho c^{2}$ to be consistent with constraints on the nuclear symmetry energy $S\simeq\rho/n-m_{n}c^{2}$, assumed to be known to be normally constrained with a mean of $32~{}{\rm MeV}$ and a standard deviation of $2~{}{\rm MeV}$ at saturation density, assumed to be 2.7 ×$\times 10^{14}~{}\rm{g~{}cm^{-3}}$ [74, 137, 138]. Figure 3 shows the marginal likelihood for different EoS realizations, plotted versus two EoS properties $R_{1.4}$ and $M_{\rm max}$. The symmetry energy constraint favors smaller radii at each fixed $M_{max}$.
Based on observations of AT2017gfo, several groups have argued that the observed electromagnetic energetics cannot be consistent with prompt collapse of the remnant to a black hole, using that qualitative constraint to infer an upper bound to the NS maximum mass; see, e.g., [54, 139, 55, 140, 55, 141] and references therein. Roughly speaking, these arguments suggest $M_{max}<2.3~{}{\rm M_{\odot}}$ [55], extremely close to the mass inferred for J0952-0607 [141].
As noted in Section II.2, the observation of AT2017gfo favors large amounts of ejecta. While these large ejecta masses can be accommodated with exceptional NS spins or mass ratios, these outcomes can also be produced with large neutron stars. Using this rationale, some groups have argued that the NS tidal deformability should be bounded below by approximately $\tilde{\Lambda}>400$ [53]. This proposed constraint on $\tilde{\Lambda}$ has comparable qualitative impact on the information obtained from GW170817 as our constraint that $q>0.9$ and $\chi_{i,z}>0$, as will be seen in Section III.2 Figure 7. For brevity, we omit further discussion of this proposal.
A preliminary analysis with the inclusion of a recently measured galactic pulsar J0437 [142] is discussed it in Appendix E.
II.6 Combining multiple measurements
We combine multiple real or synthetic observations by multiplying their likelihoods together, along with a prior $p({\cal Y})$ over the nuclear equation of state parameters ${\cal Y}$ (here ${\cal Y}=(\lambda_{1},\lambda_{2}\ldots)$ the parameters of spectral representation):
In this expression, $n$ indexes the neutron star mass radius observations and $Z_{k}$ is computed according to Eq. (6); $k$ indexes the pulsar mass measurements providing a useful constraint on the maximum neutron star mass, with ${\cal L}_{k}$ computed using one of the factors in Eq. (7); and $g$ indexes the gravitational wave observations, with the GW likelihood computed via Eq. (9). In the equation presented above, for simplicity we have not incorporated two factors superfluous to our study but essential to real analyses of the unknown neutron star populations: survey selection biases and models for the unknown neutron star population. For example, unbiased estimates of the nuclear equation of state require self-consistent estimates of the binary neutron star population [93]. In turn, self-consistent population estimates must correctly account for parameter-dependent selection effects effects [143, 144].
II.7 Hyperparameter posterior pipeline: HyperPipe
To generate the posterior distribution from ${\cal L}_{\rm net}$ and $p({\cal Y})$ we follow the RIFT iterative inference strategy HyperPipe. Starting with some proposed parameters ${\cal Y}_{\alpha}$, we evaluate ${\cal L}_{\rm net}({\cal Y}_{\alpha})\equiv{\cal L}_{net,\alpha}$; build an approximation ^$\hat{\cal L}({\cal Y})$; and use that approximation with an adaptive Monte Carlo integration program to generate both an estimated posterior distribution for ${\cal Y}$ as well as a new set of proposed parameters ${\cal Y}_{\alpha^{\prime}}$ meriting further investigation. This pipeline is able to adaptively explore the space of selected hyperparameters to resolve distinct features in the posterior distribution. The estimate is initiated in a small region of the hyperparameter space (in this case a subsection of the space of $\lambda_{1},\lambda_{2}\ldots$). We refer to the proposed parameter selection, ${\cal Y}_{\alpha}$, as the initial grid of parameters and the resultant physical properties as initial samples. Following the first iteration of evaluating likelihoods a posterior is estimated by performing Monte Carlo integration. Following this new parameters are sampled in the regions of the posterior with lower resolution in order to not exclude important finer details in addition to samples placed in exploratory regions. Following likelihood evaluations for these new points an updated posterior is evaluated by using prior and marginal likelihood information from samples of the previous iteration(s) and the current iteration, hence being better informed. Following this the process is repeated a few times where new sample points are generated and the posterior updated by Monte Carlo integration until the posterior does not deviate between iteration to iteration, resulting in a stable posterior estimate.
For this, we have extended the open-source RIFT inference package to include a general-purpose implementation of this algorithm. Our implementation not only provides a physics-neutral implementation of the RIFT algorithm, but also can interface with generic codes which compute marginal evidence factors appearing in multi-event and multi-messenger inference like Eq. (11). Moreover, though not used here as the pertinent factors are constant in this study, our implementation can include the necessary population-averaged integrals needed to account for a flexible population model and survey selection effects associated each class of measurement (i.e., the full inhomogeneous Poisson likelihood [144] or the rate-marginalized likelihood [143, 145]). Similarly, we can also easily account for multinomial leave-few-out averaged results; for example, to marginalize over including the HESS result or not, the factor ${\cal E}$ for the HESS result can be replaced by ${\cal E}p+(1-p)$ and then marginalized over confidence $p$ in that analysis.
Figure 4: Top panel: Likelihood ${\cal Z}$ versus $R_{1.4}$ deduced by comparing realizations of the $\Gamma$ spectral EoS family to the NICER mass-radius constraints of J0030 (blue) and HESS J1731 (orange) Bottom panel: Likelihood ${\cal L}$ versus $M_{max}$. The blue solid, black dashed, and orange dash-dotted lines show Eq. (7), evaluated for each known pulsar mass. The dark-red points show a direct evaluation using realizations of the $\Gamma$ spectral EoS family to the pulsar J0348, for expediency using lalsuite rather than RePrimAnd to evaluate $M_{max}$ for each EoS. The dotted green line shows the hypothesized lower bound on the maximum mass implied by prompt collapse to a black hole; see Section II.5 for discussion.
II.8 Diagnostics
@@
These diagnostics allow us to quickly propagate and interpret the effect of different prior choices about the neutron star population. For example, adopting an alternative neutron star mass distribution that is uniform out to $M_{\rm max}(\alpha)$ instead of a fixed value (such as in this study) shifts log likelihood for each individual likelihood contribution in $\ln{\cal Z}$ by a factor $-\ln\frac{M_{\rm max}-m_{min}}{2.1~{}{\rm M_{\odot}}-m_{min}}$. Because this systematic factor appears along with every mass-radius measurement, its effects stack as more measurements accumulate, increasingly penalizing configurations with exceptional $M_{\rm max}$ relative to our approach. For expected choices for $M_{\rm max}$, however, this factor is small and should have minimal impact on our posteriors.
III Results
In this section we describe inferences about the EoS, both to compare our methods against previous work and to identify challenges, along with demonstrating the impact of NS population distribution on source properties.
III.1 Inferences using galactic pulsars
@@
Figure 6: Impact of parameterization on EoS inference: Like Figure 5, but adopting the $\Upsilon$-spectral EoS, and constraining for only the PSRs J0030, J0348, J0740, J1614 and PREX symmetry energy. We also provide a benchmark analysis with the same PSRs, and symmetry energy, but includes GW constraint and uses $\Gamma$-spectral parameterization for these constraints for reference [45].
@@
We first reanalyze previously-studied events with our approach, initially using the original spectral parameterizations and priors for the assumed nuclear equation of state. To illustrate how our analysis proceeds using comparable constraints as in [45] (i.e., galactic neutron star observations and the nuclear symmetry energy), the colored scatterplot points in top left panel of Figure 5 show the marginal likelihoods versus the four model hyperparameters ($\Gamma$-spectral parameters in this case). In this example, we are employing all the galactic pulsar mass and radius measurements mentioned in Section II.3 (i.e., mass-radius information for PSR-J0030, and mass measurements for J0740, J0348, J1614), as well as the lower bounds on maximum neutron star mass from observed massive pulsars, as indicated in Figure 2. As described in Section II.7, these likelihood evaluations have been identified by HyperPipe by conducting an iterative exploration of the model space, in the process recovering the posterior distribution, whose 90% credible intervals and one-dimensional marginal distributions are also shown in the same panel of Figure 5. As expected given how broad and weak our present observational constraints on neutron star radius as illustrated in Figure 2, we find our marginal likelihoods are relatively uniform over a wide range of the parameter space, insofar as EoS compatible with those parameters exist. (Most of our nominally uniform prior volume encompasses incompatible EoS.) Next, the curves in the top-right (bottom) panels express the 90% credible interval of our inferred equation of state in $p(\rho)$ ($M(R)$). Once again, we see that our posterior 90% credible pointwise interval only differs modestly from the full range suggested by our initial assumptions, being narrower at all Mass-Radius estimates. The initial sample space is only indicative of the EoSs taken into consideration at first, as HyperPipe selects new points to explore, the posterior is free to go beyond this initial sample limits to attain the best fit. Overall, because our prior assumptions are very well adapted to contemporary theoretical preferences, the new NICER observations add relatively little new additional information over and beyond our original preferences, with one important exception highlighted below.
@@
To investigate the impact of the EoS parameterization (and prior), we also inferred the EoS with the $\Upsilon$-spectral family. Figure 6 shows a comparable analysis derived using the $\Upsilon$- spectral parameterization (e.g. Ref. 3), illustrating the inferred relationships between pressure and density, and mass and radius. In this figure we construct $p(\rho)$ posterior distributions using one-dimensional credible intervals at each density, while $M(R)$ is constructed similarly versus $M$. In sharp contrast to our fiducial analysis with the $\Gamma$-spectral parameterization, the $\Upsilon$-spectral parameterization converges to extreme EoS at low and high density. Close to our matching density with the fixed crust EoS, this $\Upsilon$-spectral posterior converges to sharp density transitions, consistent with $c_{s}\simeq 0$. At the densities $\gtrsim 10^{14.5}$ g cm-3, however, $c_{s}\to\infty$. In other words, this alternative parameterization (with broader priors) best fits galactic pulsar observations with strong low-density phase transitions and unexpected high-density behavior, so the $M(R)$ curve aligns more closely with astrophysical results, notably the NICER observation. While the EoS $p(\rho)$ and $c_{s}(\rho)$ have unexpected behavior, this model family of course produces mass-radius behavior consistent with astrophysical constraints, which are most informative in radius over a narrow mass range.
@@
$\frac{{\cal L}_{\Upsilon},{\rm max}}{{\cal L}_{\Gamma},{\rm max}}\sim 2.39~{}~{}.$ (12)
@@
Moreover, the best-fitting EoS identified by the $\Upsilon$-spectral family has qualitative differences from the EoS explored with the $\Gamma$-spectral EoS: a regime of EoSs with an early phase-transition followed by a sharp upturn such that $c_{s}\to\infty$. We attribute the differences between the two parameterizations’ results to their underlying form: given our fiducial hyperparameter priors, the $\Gamma$-spectral framework is not sufficiently flexible to produce EoSs with such sharp jumps. Our study corroborates many previous investigations on how different EoS parameterizations, priors, and phase transitions implementations can remain consistent with observations yet correspond to very different underlying conclusions about nuclear physics.
III.2 Adding information from GW sources
In order to jointly infer the nuclear EOS simultaneously using PSR and GW observations, we need the gravitational wave likelihood $L(X)$ in Eq. (9), to compute the appropriate marginal evidence ${\cal E}$ for each EOS. In this section, we first describe our reanalysis of GW170817, with which we construct the necessary marginal likelihood. We then demonstrate, using this phenomenological approach alone and without recourse to detailed EOS modeling, that the posterior for $\tilde{\Lambda}$ depends strongly on the assumed BNS population prior adopted, specifically using three concrete examples motivated in Section II.4. We discuss the correlations and features of the GW posterior that cause these prior changes to have the large effect on the $\tilde{\Lambda}$ posterior, and reflect on whether they are robust to other modeling systematics. Finally, using the likelihood provided by this analysis, we perform full-end-to-end EOS inference with the GW and PSR likelihoods together, for the different BNS population prior models discussed above. As anticipated by our simple order-of-magnitude argument using only the posterior for $\tilde{\Lambda}$, we find the BNS population priors have a substantial effect on conclusions about the nuclear EOS.
III.2.1 Reanalysis of GW170817 alone
@@
As a benchmark analysis including tidal effects, we analyze open GW data for GW170817 available from GWOSC [146], using the same PSDs provided with GWTC-1 [147, 148], over a frequency range from $23{\rm Hz}$ to $1700{\rm Hz}$ [149, 150]. By default, we adopt a low-spin prior $|\chi_{i}|<0.05$, with known sky location and source luminosity distance derived from the electromagnetically-identified host galaxy. We do not impose any constraints on the tidal deformability, allowing both $\Lambda_{i}$ to take on arbitrary values. Other prior assumptions are customary (e.g., uniform in detector-frame component masses). We include prior information about the alignment between the binary’s angular momentum direction and the line of sight, inferred from late-time radio afterglow observation [151, 152]. For simplicity and without loss of generality we assume both component spins are parallel to the orbital angular momentum direction, primarily to facilitate external constraints on the binary’s orbital inclination. Parameter inference is performed with the most recent RIFT compact binary parameter inference engine [122, 153, 154].
@@
Figure 7 shows our results for the joint posterior in mass, aligned effective inspiral spin $\chi_{\rm eff}$, and tidal deformability $\tilde{\Lambda}$, showing both the posterior (contours) and marginal likelihood (points), obtained using different prior assumptions about the BNS population. All marginal likelihoods accumulated during multiple independent analyses of GW170817, performed with different intrinsic priors but with the same data handling and priors over extrinsic parameters are shown. Some of these prior assumptions by construction limit the range of spins and $q$ allowed, so sharp edges in posteriors associated with the prior constraints are expected and natural. For the fiducial analysis (black), the marginal distributions of mass, mass ratio, spin, and $\tilde{\Lambda}$ are consistent with previously published interpretations using comparable models, settings, and data [1, 2]. For comparison and to highlight the impact of astrophysically plausible prior assumptions on source parameter inferences, this figure also shows analyses in which we require either $q>0.9$, $\chi_{i,z}>0$, or both. As the likelihood contains strong correlated features, these constraints impact the posterior distribution of tidal deformability $\tilde{\Lambda}$. As seen in the $\tilde{\Lambda}-\chi_{eff}$ 2-d posterior, when applying only the $q>0.9$ condition the strong correlation between $\tilde{\Lambda}$ and $\chi_{eff}$ requires negative $\chi_{eff}$ for small $\tilde{\Lambda}$. As a result, by further requiring that $\chi_{i}>0$, we eliminate the possibility that $\tilde{\Lambda}$ can be small (blue). I.e., when applied individually, these two constraints have modest impact. However, when imposed simultaneously, even these modest constraints produce qualitative changes in the inferred tidal deformability.
@@
Figure 7: GW inference Benchmark inference of properties of GW170817 using RIFT and IMRPhenomPv2_NRTidalv2. Color scale shows the inferred marginal likelihood, increasing from blue to red over a range of 15 in log-likelihood (with gray indicating values more than $15$ below the peak); 2d contours show the 90% credible intervals; and one-dimensional distributions show the one-dimensional marginal posterior distributions. Parameters shown are the detector-frame chirp mass (denoted here by ${\cal M}_{c}$), mass ratio $q=m_{2}/m_{1}$, aligned effective inspiral spin $\chi_{\rm eff}$, and tidal deformability $\tilde{\Lambda}$. Results for our fiducial analysis (in black) shown here are very similar to published results available through GWOSC [1, 2]. Colors indicate results derived using fiducial priors (black); $q>0.8$ (red); $\chi_{i,z}>0$ (green); and both (blue). The magenta curve shows results derived with a spin prior extending up to $|\chi_{i}|=0.4$.
@@
Figures 7 and 8 also illustrate inferences derived using even more extreme prior assumptions about the pertinent neutron star spin and mass ratio. For example, for the magenta curve provided in Figure 7, the analysis merely allowed for larger possible values of the NS maximum spin. Relaxing the upper limit on $\chi_{i}$ allows a lower mass for $m_{2}$ to remain consistent with the data [2], while increasing $\chi_{i}$ only incrementally. Though this single changed assumption allows for more extreme masses, in isolation this one relaxed assumption doesn’t change our qualitative interpretation; notably, the distribution of $\tilde{\Lambda}$ is nearly unchanged. By contrast, Figure 8 shows inferences derived assuming $m_{2}/m_{1}<0.6$, both adopting fiducial spin magnitudes (black) and larger NS spins (red). Notably, the likelihood includes comparable values for a wide range of mass ratios, including masses $m_{2}<1~{}{\rm M_{\odot}}$. Contrasting the likelihood (points) to the two posterior marginals (black and red) shows how strongly the arbitrary upper limit in our fiducial spin prior limits the mass ratio distribution. Both extreme-$q$ analyses nonetheless produce comparable $\tilde{\Lambda}$ posterior distributions. Due to correlations between mass ratio and $\tilde{\Lambda}$, the constraint $q<0.6$ causes the posterior distribution of $\tilde{\Lambda}$ to systematically decrease. In Appendix C, we show that the median $\tilde{\Lambda}$ steadily decreases versus an imposed maximum mass ratio $q$.
@@
These GW-only reanalyses suggest that hidden astrophysical priors change $\tilde{\Lambda}$ by roughly a factor of 2, increasing the median $\tilde{\Lambda}$ (if we require comparable mass and positive aligned spins) or decreasing it (if we require $q<0.6$) respectively. However, because of the exponent relating $R_{1.4}$ and $\tilde{\Lambda}$, see Eq. (10), these factors of 2 propagate to a roughly 10% change in the inferred $R_{1.4}$, comparable to the statistical measurement errors usually presented for many of the individual measurements described earlier in Section II.3. Figure 9 shows the cumulative distribution of $R_{1.4}$ derived from this expression and three GW-only analyses: our fiducial analysis (blue); the comparable mass positive spin analysis (orange); and the high mass ratio analysis (green). Note, the bimodality in the probability distributions of the fiducial and the comparable-mass positive spin analyses arise due to bimodality in $\tilde{\Lambda}$ seen in Fig. 7.

### v3

@@ 2.9 Hyperparameter posterior pipeline: HyperPipe
[…]
2.10 Diagnostics
@@
These diagnostics allow us to quickly propagate and interpret the effect of different prior choices about the neutron star population. For example, adopting an alternative neutron star mass distribution that is uniform out to $M_{\rm max}(\alpha)$ instead of a fixed value (such as in this study) shifts log likelihood for each individual likelihood contribution in $\ln{\cal Z}$ by a factor $\ln\frac{M_{\rm max}-m_{min}}{2.1\penalty 10000\ {\rm M_{\odot}}-m_{min}}$. Because this systematic factor appears along with every mass-radius measurement, its effects stack as more measurements accumulate, increasingly penalizing configurations with exceptional $M_{\rm max}$ relative to our approach. For expected choices for $M_{\rm max}$, however, this factor is small and should have minimal impact on our posteriors.
3 Results
In this section we describe inferences about the EoS. In Section 3.1 we first demonstrate our method using only galactic pulsars and the nuclear symmetry energy. We show how our HyperPipe implementation, like RIFT, provides interpretable net marginal likelihoods, allowing us to visually assess the posterior and our coverage similar to typical RIFT analyses. Using the concrete example of the two spectral EoS parameterizations described earlier ($\Gamma$ and $\Upsilon$), we show that different EoS model families can produce qualitatively different results. In Section 3.2, we describe how we incorporate information from GW170817 into EoS inference using different assumptions about the pertinent binary NS population. Specfically, in Section 3.2.1 we describe EoS-agnostic RIFT source parameter inference of GW170817 conditioned on several different alternatives. population. We demonstrate that even before explicitly conditoning on an EoS model family, these population assumptions have substantial impact on the range of $\tilde{\Lambda}$ with significant support in each scenario. Then, in Section 3.2.2 we present EoS inference conditioned on our fiducial assumptions as well as with these alternative binary NS population assumptions. We find the choice of NS population assumptions will significantly change the inferred radius of a typical NS. In Section 3.3, for context and a sense of scale, we review for comparison the extent to which additional ad hoc constraints on the NS maximum mass can further inform the EoS. Similarly and last, in Section 3.4 we discuss the marginal impact that the HESS J1731 source has on our overall inferences, compared to alternative binary NS assumptions for GW170817.
3.1 Inferences using galactic pulsars
@@
Figure 6: Impact of parameterization on EoS inference: Like Figure 5, but adopting the $\Upsilon$-spectral EoS, and constraining for only the PSRs J0030, J0348, J0740, J1614 and PREX symmetry energy. We also provide a comparison analysis which adopted the same PSRs, and symmetry energy, but includes GW constraint and uses $\Gamma$-spectral parameterization for these constraints for reference (Miller:2021qha).
@@
We first reanalyze previously-studied events with our approach, initially using the original spectral parameterizations and priors for the assumed nuclear equation of state. To illustrate how our analysis proceeds using comparable constraints as in (Miller:2021qha) (i.e., galactic neutron star observations and the nuclear symmetry energy), the colored scatterplot points in top left panel of Figure 5 show the marginal likelihoods versus the four model hyperparameters ($\Gamma$-spectral parameters in this case). In this example, we are employing all the galactic pulsar mass and radius measurements mentioned in Section 2.4 (i.e., mass-radius information for PSR-J0030, and mass measurements for J0740, J0348, J1614), as well as the lower bounds on maximum neutron star mass from observed massive pulsars, as indicated in Figure 1. As described in Section 2.9, these likelihood evaluations have been identified by HyperPipe by conducting an iterative exploration of the model space, in the process recovering the posterior distribution, whose 90% credible intervals and one-dimensional marginal distributions are also shown in the same panel of Figure 5. As expected given how broad and weak our present observational constraints on neutron star radius as illustrated in Figure 1, we find our marginal likelihoods are relatively uniform over a wide range of the parameter space, insofar as EoS compatible with those parameters exist. (Most of our nominally uniform prior volume encompasses incompatible EoS.) Next, the curves in the top-right (bottom) panels express the 90% credible interval of our inferred equation of state in $p(\rho)$ ($M(R)$). Once again, we see that our posterior 90% credible pointwise interval only differs modestly from the full range suggested by our initial assumptions, being narrower at all Mass-Radius estimates. The initial sample space is only indicative of the EoSs taken into consideration at first, as HyperPipe selects new points to explore, the posterior is free to go beyond this initial sample limits to attain the best fit. Overall, because our prior assumptions are very well adapted to contemporary theoretical preferences, the new NICER observations add relatively little new additional information over and beyond our original preferences, with one important exception highlighted below.
@@
To investigate the impact of the EoS parameterization (and prior), we also inferred the EoS with the $\Upsilon$-spectral family. Figure 6 shows a comparable analysis derived using the $\Upsilon$- spectral parameterization (e.g. Eq. 10), illustrating the inferred relationships between pressure and density, and mass and radius. In this figure we construct $p(\rho)$ posterior distributions using one-dimensional credible intervals at each density, while $M(R)$ is constructed similarly versus $M$. In sharp contrast to our fiducial analysis with the $\Gamma$-spectral parameterization, the $\Upsilon$-spectral parameterization converges to extreme EoS at low and high density. Close to our matching density with the fixed crust EoS, this $\Upsilon$-spectral posterior converges to sharp density transitions, consistent with $c_{s}\simeq 0$. At the densities $\gtrsim 10^{14.5}$ g cm-3, however, $c_{s}\to 1$. In other words, this alternative parameterization (with broader priors) best fits galactic pulsar observations with strong low-density phase transitions and unexpected high-density behavior, so the $M(R)$ curve aligns more closely with astrophysical results, notably the NICER observation. While the EoS $p(\rho)$ and $c_{s}(\rho)$ have unexpected behavior, this model family of course produces mass-radius behavior consistent with astrophysical constraints, which are most informative in radius over a narrow mass range.
@@
$\frac{{\cal L}_{\Upsilon},{\rm max}}{{\cal L}_{\Gamma},{\rm max}}\sim 2.39\penalty 10000\ \penalty 10000\ .$ (12)
@@
Moreover, the best-fitting EoS identified by the $\Upsilon$-spectral family has qualitative differences from the EoS explored with the $\Gamma$-spectral EoS: a regime of EoSs with an early phase-transition followed by a sharp upturn such that $c_{s}\to 1$. We attribute the differences between the two parameterizations’ results to their underlying form: given our fiducial hyperparameter priors, the $\Gamma$-spectral framework is not sufficiently flexible to produce EoSs with such sharp jumps. Our study corroborates many previous investigations on how different EoS parameterizations, priors, and phase transitions implementations can remain consistent with observations yet correspond to very different underlying conclusions about nuclear physics.
3.2 Adding information from GW sources
[…]
3.2.1 Reanalysis of GW170817 alone
@@
As a benchmark analysis including tidal effects, we analyze open GW data for GW170817 available from GWOSC (ligo-O1O2-opendata), using the same PSDs provided with GWTC-1 (LIGO-O2-Catalog; LIGO-O2-Catalog-PSDRelease), over a frequency range from $23{\rm Hz}$ to $1700{\rm Hz}$ (TheLIGOScientific:2014jea; TheVirgo:2014hva). By default, we adopt a low-spin prior $|\chi_{i}|<0.05$, with known sky location and source luminosity distance derived from the electromagnetically-identified host galaxy. We do not impose any constraints on the tidal deformability, allowing both $\Lambda_{i}$ to take on arbitrary values. Other prior assumptions are customary (e.g., uniform in detector-frame component masses). We include prior information about the alignment between the binary’s angular momentum direction and the line of sight, inferred from late-time radio afterglow observation (2019NatAs...3..940H; 2018Natur.554..207M). For simplicity and without loss of generality we assume both component spins are parallel to the orbital angular momentum direction, primarily to facilitate external constraints on the binary’s orbital inclination. Parameter inference is performed with the most recent RIFT compact binary parameter inference engine (gwastro-PENR-RIFT; gwastro-PENR-RIFT-GPU; PhysRevD.107.024040).
@@
Figure 7 shows our results for the joint posterior in mass, aligned effective inspiral spin $\chi_{\rm eff}$, and tidal deformability $\tilde{\Lambda}$, showing both the posterior (contours) and marginal likelihood (points), obtained using different prior assumptions about the BNS population. All marginal likelihoods accumulated during multiple independent analyses of GW170817, performed with different intrinsic priors but with the same data handling and priors over extrinsic parameters are shown. Some of these prior assumptions by construction limit the range of spins and $q$ allowed, so sharp edges in posteriors associated with the prior constraints are expected and natural. For the fiducial analysis (black), the marginal distributions of mass, mass ratio, spin, and $\tilde{\Lambda}$ are consistent with previously published interpretations using comparable models, settings, and data (LIGO-GW170817-bns; LIGO-GW170817-SourceProperties). For comparison and to highlight the impact of astrophysically plausible prior assumptions on source parameter inferences, this figure also shows analyses in which we require either $q>0.9$, $\chi_{i,z}>0$, or both. As the likelihood contains strong correlated features, these constraints impact the posterior distribution of tidal deformability $\tilde{\Lambda}$. As seen in the $\tilde{\Lambda}-\chi_{\rm eff}$ 2-d posterior, when applying only the $q>0.9$ condition the strong correlation between $\tilde{\Lambda}$ and $\chi_{\rm eff}$ requires negative $\chi_{\rm eff}$ for small $\tilde{\Lambda}$. As a result, by further requiring that $\chi_{i}>0$, we eliminate the possibility that $\tilde{\Lambda}$ can be small (blue). I.e., when applied individually, these two constraints have modest impact. However, when imposed simultaneously, even these modest constraints produce qualitative changes in the inferred tidal deformability.
@@
Figure 7: GW inference Benchmark inference of properties of GW170817 using RIFT and IMRPhenomPv2_NRTidalv2. Color scale shows the inferred marginal likelihood, increasing from blue to red over a range of 15 in log-likelihood (with gray indicating values more than 1515 below the peak); 2d contours show the 90% credible intervals; and one-dimensional distributions show the one-dimensional marginal posterior distributions. Parameters shown are the detector-frame chirp mass (denoted here by ${\cal M}_{c}$), mass ratio $q=m_{2}/m_{1}$, aligned effective inspiral spin $\chi_{\rm eff}$, and tidal deformability $\tilde{\Lambda}$. Results for our fiducial analysis (in black) shown here are very similar to published results available through GWOSC (LIGO-GW170817-bns; LIGO-GW170817-SourceProperties). Colors indicate results derived using fiducial priors (black); $q>0.8$ (red); $\chi_{i,z}>0$ (green); and both (blue). The magenta curve shows results derived with a spin prior extending up to $|\chi_{i}|=0.4$.
@@
Figures 7 and 8 also illustrate inferences derived using even more extreme prior assumptions about the pertinent neutron star spin and mass ratio. For example, for the magenta curve provided in Figure 7, the analysis merely allowed for larger possible values of the NS maximum spin. Relaxing the upper limit on $\chi_{i}$ allows a lower mass for $m_{2}$ to remain consistent with the data (LIGO-GW170817-SourceProperties), while increasing $\chi_{i}$ only incrementally. Though this single changed assumption allows for more extreme masses, in isolation this one relaxed assumption doesn’t change our qualitative interpretation; notably, the distribution of $\tilde{\Lambda}$ is nearly unchanged. By contrast, Figure 8 shows inferences derived assuming $m_{2}/m_{1}<0.6$, both adopting fiducial spin magnitudes (black) and larger NS spins (red). Notably, the likelihood includes comparable values for a wide range of mass ratios, including masses $m_{2}<1\penalty 10000\ {\rm M_{\odot}}$. Contrasting the likelihood (points) to the two posterior marginals (black and red) shows how strongly the arbitrary upper limit in our fiducial spin prior limits the mass ratio distribution. Both extreme-$q$ analyses nonetheless produce comparable $\tilde{\Lambda}$ posterior distributions. Due to correlations between mass ratio and $\tilde{\Lambda}$, the constraint $q<0.6$ causes the posterior distribution of $\tilde{\Lambda}$ to systematically decrease. In Appendix C, we show that the median $\tilde{\Lambda}$ steadily decreases versus an imposed maximum mass ratio $q$.
@@
These GW-only reanalyses suggest that astrophysical priors change $\tilde{\Lambda}$ by roughly a factor of 2, increasing the median $\tilde{\Lambda}$ (if we require comparable mass and positive aligned spins) or decreasing it (if we require $q<0.6$) respectively. However, because of the exponent relating $R_{1.4}$ and $\tilde{\Lambda}$, see Eq. (7), these factors of 2 propagate to a roughly 10% change in the inferred $R_{1.4}$, comparable to the statistical measurement errors usually presented for many of the individual measurements described earlier in Section 2.4. Figure 9 shows the cumulative distribution of $R_{1.4}$ derived from this expression and three GW-only analyses: our fiducial analysis (blue); the comparable mass positive spin analysis (orange); and the high mass ratio analysis (green). Note, the bimodality in the probability distributions of the fiducial and the comparable-mass positive spin analyses arise due to bimodality in $\tilde{\Lambda}$ seen in Fig. 7.